# Degrees of Freedom for Radiating Systems

### v1

@@ VI Inverse source problems
which resembles the communication model in Sec III. In inverse source problems, the source current $\mathbf{I}$ is estimated from the observed field $\mathbf{f}$. The system (29) is not invertible and it is common to reformulate the system as an optimization problem or to use an SVD for a regularized solution [36].
@@ VII Conclusions
We present a method for determining the number of degrees of freedom (NDoF) for a radiating system based on radiation modes. Asymptotically, the NDoF approaches the total shadow area measured in squared wavelengths. For convex regions, the NDoF matches the results derived from Weyl’s law. The findings extend the known expressions to arbitrary shapes, encompassing non-convex and non-connected regions. Numerical results complementing the theoretical framework and comparing them with the shadow area are provided. The derivation of the NDoF is linked to upper limits on the capacity (spectral efficiency) for a communication channel between an object and the far field, as well as to inverse source problems with measurements surrounding an object.
@@ Appendix A Radiation modes
$\varrho_{n}=\frac{k^{2}a^{2}\eta_{0}}{R_{\mathrm{s}}}(\mathrm{R}_{\tau,l}^{(1)})^{2}$ (36)

### v2

@@ VI Inverse source problem
which resembles the communication model in Sec III. In inverse source problems, the source current $\mathbf{I}$ is estimated from the observed field $\mathbf{f}$. The system (27) is not invertible, and it is common to reformulate the system as an optimization problem or to use an SVD for a regularized solution [42].
@@ VII Conclusions
We present a method for determining the NDoF for a radiating system based on radiation modes. Asymptotically, the NDoF approaches the total shadow area measured in squared wavelengths. For convex regions, the NDoF matches the results derived from Weyl’s law. The findings extend the known expressions to arbitrary shapes, encompassing non-convex and non-connected regions. Numerical results complementing the theoretical framework and comparing them with the shadow area are provided. The derivation of the NDoF is linked to upper limits on the capacity (spectral efficiency) for a communication channel between an object and the far field, as well as to inverse source problems with measurements surrounding an object.
@@ Appendix A Radiation modes
$\varrho_{n}=\frac{k^{2}a^{2}\eta_{0}}{R_{\mathrm{s}}}(\mathrm{R}_{\tau,l}^{(1)})^{2},$ (34)
@@
Appendix B Directional and polarization averages
Expressing the solution (15) in spherical waves with expansion coefficients collected in a column matrix $\mathbf{a}$ using $\mathbf{F}=\mathbf{U}^{\text{T}}\mathbf{a}/(4\pi)$ [27] yields
$\max A_{\mathrm{eff}}=\lambda^{2}\mathbf{F}^{\text{H}}\mathbf{R}^{-1}\mathbf{F}=\frac{\lambda^{2}}{16\pi^{2}}\mathbf{a}^{\text{H}}\mathbf{U}\mathbf{R}^{-1}\mathbf{U}^{\text{T}}\mathbf{a}.$ (35)
Diagonalizing (35) using radiation modes (11) induces a change of variables $\tilde{\mathbf{a}}=\mathbf{Q}\mathbf{a}$ with a unitary matrix $\mathbf{Q}$ having elements $Q_{nm}$
with $\sum_{m}Q^{2}_{nm}=1$. Taking the average (16) of $\max A_{\mathrm{eff}}$ and using the plane wave expansion $a_{n}=4\pi\mathrm{j}^{\tau-1-l}\hat{\boldsymbol{e}}\cdot\boldsymbol{Y}_{n}(\hat{\boldsymbol{k}})$ [23] with spherical harmonics $\boldsymbol{Y}_{n}$ and $\hat{\boldsymbol{e}}\cdot\hat{\boldsymbol{k}}=0$ results in
and using that $\hat{\boldsymbol{k}}\cdot\boldsymbol{Y}(\hat{\boldsymbol{k}})=0$ extends (38) to an identity dyadic. The same relation (37) holds for the orthogonal combination $\langle|\tilde{a}_{n}|^{2}\rangle=\sum Q^{2}_{mn}\langle|a_{n}|^{2}\rangle=2\pi$ giving the simple identity (17) for the average maximal effective area expressed in radiation modes.